Determine the amount of sales (units) that would be necessary under Break-Even Sales Under Present and...

  1. Determine the amount of sales (units) that would be necessary under

    Break-Even Sales Under Present and Proposed Conditions

    Darby Company, operating at full capacity, sold 74,250 units at a price of $51 per unit during the current year. Its income statement for the current year is as follows:

    Sales $3,786,750
    Cost of goods sold 1,870,000
    Gross profit $1,916,750
    Expenses:
    Selling expenses $935,000
    Administrative expenses 935,000
    Total expenses 1,870,000
    Income from operations $46,750

    The division of costs between fixed and variable is as follows:

    Variable Fixed
    Cost of goods sold 70% 30%
    Selling expenses 75% 25%
    Administrative expenses 50% 50%

    Management is considering a plant expansion program that will permit an increase of $306,000 in yearly sales. The expansion will increase fixed costs by $30,600, but will not affect the relationship between sales and variable costs.

    Required:

    1. Determine the total variable costs and the total fixed costs for the current year. Enter the final answers rounded to the nearest dollar.

    Total variable costs $
    Total fixed costs $

    2. Determine (a) the unit variable cost and (b) the unit contribution margin for the current year. Enter the final answers rounded to two decimal places.

    Unit variable cost $
    Unit contribution margin $

    3. Compute the break-even sales (units) for the current year. Enter the final answers rounded to the nearest whole number.
    units

    4. Compute the break-even sales (units) under the proposed program for the following year. Enter the final answers rounded to the nearest whole number.
    units

    5. Determine the amount of sales (units) that would be necessary under the proposed program to realize the $46,750 of income from operations that was earned in the current year. Enter the final answers rounded to the nearest whole number.
    units

    6. Determine the maximum income from operations possible with the expanded plant. Enter the final answer rounded to the nearest dollar.
    $

    7. If the proposal is accepted and sales remain at the current level, what will the income or loss from operations be for the following year? Enter the final answer rounded to the nearest dollar.
    $  

    8. Based on the data given, would you recommend accepting the proposal?

    1. In favor of the proposal because of the reduction in break-even point.
    2. In favor of the proposal because of the possibility of increasing income from operations.
    3. In favor of the proposal because of the increase in break-even point.
    4. Reject the proposal because if future sales remain at the current level, the income from operations will increase.
    5. Reject the proposal because the sales necessary to maintain the current income from operations would be below the current year sales.

    Choose the correct answer.
Answer #1

as per HomeworkLib policy i have answered first 4 question for rest of question please raise new request.. please do provide rating..

Given that-
Sales $3,786,750
Cost of goods sold 1,870,000
Gross profit $1,916,750
Expenses:
Selling expenses $935,000
Administrative expenses 935,000
Total expenses 1,870,000
Income from operations $46,750
Variable Fixed
Cost of goods sold 70% 30%
Selling expenses 75% 25%
Administrative expenses 50% 50%
Answer 1)
Variable Fixed Variable Fixed
Cost of goods sold 1,870,000 70% 30% 1309000 561000
Selling expenses $935,000 75% 25% 701250 233750
Administrative expenses 935,000 50% 50% 467500 467500
Total 3,740,000 2477750 1262250
Total variable costs 2477750
Total fixed costs 1262250
Answer 2)
Computation of unit variable cost
i ii=i/74250
Total variable cost Unit variable cost
Cost of goods sold 1309000 $             17.63
Selling expenses 701250 $               9.44
Administrative expenses 467500 $               6.30
Total 2477750 $             33.37
Computation of unit Contribution = Sales per unit - Variable cost per unit
Sales per unit $        51.00
Variable cost per unit $        33.37
Contribution per unit $        17.63
Unit variable cost $        33.37
Unit contribution margin $        17.63
Answer 3) Compute the break-even sales (units) for the current year
Break even sales = Total fixed cost / Contribution per unit
i Total fixed cost = 1262250
ii Contribution per unit = $        17.63
iii=i/ii Break even unit =          71,598
Answer 4) Compute the break-even sales (units) under the proposed program for the following year
i Total fixed cost = 1262250+30,600 $ 1,292,850.00
ii Contribution per unit = $             17.63
iii=i/ii Break even unit =               73,334